OpenStudy (anonymous):

@satellite73 @satellite73 @satellite73 MEDALS AND FAN AND TESTAMONIALS

OpenStudy (anonymous):

Which is the ordered pair for the translated point C'? C(0, –4) translated left 2, down 6 A. (–2, –10) B. (–2, 2) C. (2, 2) D. (2, 10)

OpenStudy (anonymous):

@surjithayer help plz

OpenStudy (anonymous):

X=x-2 Y=y-6 put x=0,y=-4

OpenStudy (anonymous):

D

OpenStudy (anonymous):

no X=0-2=-2 find Y

OpenStudy (anonymous):

so C and D are out of the question right ?

OpenStudy (anonymous):

process of elimination

OpenStudy (anonymous):

Y=-4-6=?

OpenStudy (anonymous):

-10

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

so A

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Which is the ordered pair for the translated point G'? Point G translated right 0, up 5 A. (–4, 1) B. (–4, 9) C. (–4, 4) D. (1, 4) http://static.k12.com/calms_media/media/1419500_1420000/1419560/1/6e8d179d12e56ca8b6089fc1dff30e32472eed83/FGA_130921_1614041.jpg

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@alexwee123

OpenStudy (anonymous):

X=x+a,if translated right ,add a ,if left subtract a Y=y+b,if translated up add b ,if down subtract b.